Prove that, if a, b, c are non-zero and



(a x b) x c = a x (b x c)



then either



(i) b is perpendicular to both a and c, or



(ii) a and c are parallel or anti-parallel




Note: "x" is cross product


1
Expert's answer
2022-01-11T00:58:34-0500
"\vec a \times (\vec b\times \vec c)=\vec b(\vec a\cdot \vec c)-\vec c(\vec a\cdot \vec b)"

"(\vec a\times \vec b)\times \vec c=-\vec c \times(\vec a\times \vec b) =-\vec a(\vec c\cdot \vec b)+\vec b(\vec c\cdot \vec a)"

Given "(\vec a\times \vec b)\times \vec c=\vec a \times (\vec b\times \vec c)."

Then


"-\vec a(\vec c\cdot \vec b)+\vec b(\vec c\cdot \vec a)=\vec b(\vec a\cdot \vec c)-\vec c(\vec a\cdot \vec b)"

Use that "\vec b(\vec c\cdot \vec a)=\vec b(\vec a\cdot \vec c), \forall\vec a,\vec b,\vec c."

Then


"-\vec a(\vec c\cdot \vec b)=-\vec c(\vec a\cdot \vec b)"




"\vec a(\vec c\cdot \vec b)=\vec c(\vec a\cdot \vec b)"

(i)

Suppose that the nonzero vectors "\vec a" and "\vec c" are not collinear.

Now suppose that "\vec b" is not perpendicular to "\vec a" . It means that "(\vec a\cdot \vec b)\not=0."

Then "(\vec c\cdot \vec b)\not=0" and

"\vec a=\dfrac{(\vec a\cdot \vec b)}{(\vec c\cdot \vec b)}\vec c"

If two nonzero vectors "\vec a" and "\vec c" are not collinear, then we cannot find the number "k\not=0" such that the "\vec a=k\vec c." Hence we have a contradiction.

Therefore if vectors "\vec a" and "\vec c" are not collinear then


"\vec c\cdot \vec b=0\ and \ \vec a\cdot \vec b=0"

These mean that "\vec b" is perpendicular to both "\vec a" and "\vec c."


(ii)

"\vec c\cdot \vec b\not=0\ or \ \vec a\cdot \vec b\not=0"


Zero vector "\vec 0=0\cdot\vec v" is considered to be parallel to every other vector "\vec v."

Then the vectors "\vec a" and "\vec b" are collinear. These mean that "\vec a" and "\vec c" are parallel or anti-parallel.